# PRISM2: Unlocking Multi-Modal General Pathology AI with Clinical Dialogue

### v1

@@ A.1 Evaluation tasks and data
Biomarker prediction is an in-house dataset containing 37,633 WSIs gathered from 33,949 specimens. The ground truth labels span 10 different biomarkers: Prostate AR, Ovarian FGA, Esophagogastric Her2, Colorectal MSI, Lung EGFR, Melanoma BRAF, Bladder FGFR, Breast CDH1, Endometrial PTEN, and Breast PIK3CA. These labels were originally identified by MSK-IMPACT [cheng2015-msk-impact], a targeted test for genetic mutations. The dataset is split into train, tune, and test subsets of size 22,634, 5,606, and 5,709 specimens respectively.
A.2 Supplementary results
Model Method PanCancer PanCancer Rare Breast Subtyping GI Suite
TITAN Contrastive 0.805 (0.794, 0.815) 0.767 (0.750, 0.785) - -
PRISM Contrastive 0.896 (0.888, 0.903) 0.851 (0.835, 0.865) - -
PRISM2 Contrastive 0.805 (0.794, 0.815) 0.804 (0.787, 0.820) - -
Dialogue (YN) 0.933 (0.926, 0.939) 0.929 (0.920, 0.938) 0.958 (0.948, 0.967) 0.932 (0.921, 0.942)
Dialogue (DYN) 0.943 (0.938, 0.949) 0.943 (0.934, 0.951) 0.956 (0.943, 0.967) 0.929 (0.917, 0.941)
(a) ROC AUC for binary classification tasks based on MSK slides.
Model Method BRCA NSCLC RCC
TITAN Contrastive 0.953 (0.935, 0.969) 0.985 (0.978, 0.991) 0.978 (0.970, 0.985)
PRISM Contrastive 0.935 (0.911, 0.956) 0.969 (0.960, 0.978) 0.919 (0.905, 0.933)
PRISM2 Contrastive 0.949 (0.926, 0.967) 0.980 (0.973, 0.986) 0.981 (0.975, 0.987)
Dialogue (OPN) 0.963 (0.946, 0.977) 0.977 (0.968, 0.985) 0.978 (0.970, 0.984)
(b) ROC AUC for multi-class classification tasks based on TCGA slides.
Table A.2.1: Zero-shot ROC AUC for various downstream classification tasks using dialogue completions. 95% confidence intervals computed with 1000 iterations of bootstrapping are shown in parentheses. The best result per task (p $<$ 0.05) is in bold.
PRISM2 Dialogue ROC AUC Balanced Accuracy
YN DYN YN DYN
Invasive ductal carcinoma (IDC) 0.984 (0.978, 0.989) 0.979 (0.972, 0.985) 0.954 (0.944, 0.964) 0.946 (0.934, 0.957)
Invasive lobular carcinoma (ILC) 0.983 (0.976, 0.989) 0.982 (0.961, 0.994) 0.863 (0.811, 0.917) 0.872 (0.821, 0.920)
Ductal carcinoma in situ (DCIS) 0.970 (0.963, 0.976) 0.963 (0.955, 0.971) 0.913 (0.900, 0.926) 0.922 (0.910, 0.933)
Lobular carcinoma in situ (LCIS) 0.960 (0.949, 0.969) 0.964 (0.951, 0.974) 0.796 (0.765, 0.826) 0.778 (0.746, 0.808)
Atypical ductal hyperplasia (ADH) 0.896 (0.877, 0.914) 0.899 (0.882, 0.915) 0.786 (0.752, 0.817) 0.673 (0.638, 0.706)
Atypical lobular hyperplasia (ALH) 0.927 (0.909, 0.944) 0.936 (0.918, 0.952) 0.811 (0.779, 0.843) 0.717 (0.684, 0.749)
Invasive carcinoma 0.979 (0.974, 0.984) 0.965 (0.956, 0.973) 0.936 (0.925, 0.946) 0.940 (0.929, 0.950)
Benign or unremarkable 0.965 (0.957, 0.973) 0.960 (0.951, 0.968) 0.930 (0.920, 0.940) 0.924 (0.913, 0.934)
Average 0.958 (0.948, 0.967) 0.956 (0.943, 0.967) 0.874 (0.849, 0.897) 0.847 (0.822, 0.870)
Table A.2.2: Zero-shot AUC and balanced accuracy of detection tasks using dialogue completions in breast tissue, spanning cancers (IDC, ILC) and precursors to cancer (DCIS, LCIS, ADH, ALH). 95% confidence intervals computed with 1000 iterations of bootstrapping are shown in parentheses. The best result per task (p $<$ 0.05) is in bold.
PRISM2 Dialogue ROC AUC Balanced Accuracy
YN DYN YN DYN
CANCER
Adenocarcinoma 0.992 (0.990, 0.994) 0.992 (0.990, 0.994) 0.966 (0.960, 0.971) 0.936 (0.926, 0.945)
Lymphoma 0.889 (0.865, 0.910) 0.906 (0.881, 0.929) 0.663 (0.629, 0.699) 0.784 (0.748, 0.820)
Metastatic Tumor 0.973 (0.958, 0.987) 0.966 (0.948, 0.982) 0.894 (0.864, 0.923) 0.829 (0.792, 0.863)
Neuroendocrine Tumor 0.964 (0.950, 0.976) 0.951 (0.932, 0.967) 0.883 (0.856, 0.908) 0.873 (0.844, 0.901)
Squamous Cell Carcinoma 0.995 (0.989, 0.998) 0.994 (0.987, 0.998) 0.949 (0.932, 0.967) 0.923 (0.899, 0.946)
Tumor Differentiation 0.991 (0.988, 0.993) 0.993 (0.990, 0.995) 0.954 (0.946, 0.961) 0.942 (0.933, 0.951)
PRECURSOR OR BENIGN
Dysplasia 0.802 (0.783, 0.820) 0.781 (0.762, 0.800) 0.713 (0.695, 0.729) 0.684 (0.664, 0.701)
Polyp 0.945 (0.940, 0.949) 0.957 (0.953, 0.961) 0.790 (0.782, 0.799) 0.849 (0.841, 0.857)
No Significant Abnormalities 0.866 (0.860, 0.872) 0.906 (0.901, 0.911) 0.599 (0.595, 0.604) 0.560 (0.556, 0.564)
INFLAMMATORY CONDITION
H. Pylori Helicobacter 0.932 (0.908, 0.953) 0.892 (0.859, 0.921) 0.536 (0.518, 0.556) 0.823 (0.788, 0.859)
Barretts Esophagus 0.972 (0.964, 0.979) 0.946 (0.935, 0.956) 0.872 (0.857, 0.886) 0.831 (0.815, 0.848)
Other Inflammatory Condition 0.864 (0.856, 0.872) 0.870 (0.862, 0.878) 0.748 (0.739, 0.758) 0.750 (0.740, 0.758)
Average 0.932 (0.921, 0.942) 0.929 (0.917, 0.941) 0.797 (0.781, 0.813) 0.815 (0.796, 0.834)
Table A.2.3: Zero-shot AUC and balanced accuracy of detection tasks using dialogue completions in the gastrointestinal (GI) tract, spanning cancers, precursors to cancer, inflammatory conditions, and benign conditions. 95% confidence intervals computed with 1000 iterations of bootstrapping are shown in parentheses. The best result per task (p $<$ 0.05) is in bold.
Linear probing PRISM2 PRISM2 PRISM TITAN
Diagnostic Base
Invasive ductal carcinoma (IDC) 0.993 (0.990, 0.995) 0.985 (0.978, 0.991) 0.988 (0.982, 0.992) 0.954 (0.944, 0.963)
Invasive lobular carcinoma (ILC) 0.993 (0.988, 0.996) 0.987 (0.979, 0.994) 0.991 (0.979, 0.997) 0.958 (0.934, 0.977)
Ductal carcinoma in situ (DCIS) 0.985 (0.980, 0.989) 0.983 (0.978, 0.987) 0.982 (0.976, 0.986) 0.938 (0.927, 0.946)
Lobular carcinoma in situ (LCIS) 0.973 (0.962, 0.981) 0.973 (0.961, 0.981) 0.971 (0.961, 0.979) 0.893 (0.870, 0.914)
Atypical ductal hyperplasia (ADH) 0.947 (0.931, 0.961) 0.943 (0.927, 0.956) 0.944 (0.928, 0.958) 0.865 (0.836, 0.890)
Atypical lobular hyperplasia (ALH) 0.972 (0.956, 0.984) 0.965 (0.947, 0.979) 0.960 (0.943, 0.976) 0.887 (0.860, 0.910)
Average 0.979 (0.976, 0.982) 0.976 (0.972, 0.979) 0.975 (0.971, 0.978) 0.923 (0.916, 0.929)
Table A.2.4: Linear probing AUC of detection tasks in breast tissue, spanning cancers (IDC, ILC and precursors to cancer (DCIS, LCIS, ADH, ALH). 95% confidence intervals computed with 1000 iterations of bootstrapping are shown in parentheses. The best result per task (p $<$ 0.05) is in bold.
Linear probing PRISM2 PRISM2 PRISM TITAN
Diagnostic Base
CANCER
Adenocarcinoma 0.995 (0.993, 0.997) 0.995 (0.994, 0.997) 0.994 (0.991, 0.996) 0.993 (0.991, 0.995)
Lymphoma 0.980 (0.970, 0.989) 0.979 (0.967, 0.990) 0.971 (0.957, 0.983) 0.970 (0.955, 0.982)
Metastatic Tumor 0.989 (0.978, 0.997) 0.987 (0.977, 0.995) 0.978 (0.964, 0.990) 0.971 (0.956, 0.985)
Neuroendocrine Tumor 0.983 (0.970, 0.992) 0.987 (0.978, 0.993) 0.981 (0.968, 0.990) 0.977 (0.962, 0.989)
Squamous Cell Carcinoma 0.997 (0.994, 0.999) 0.995 (0.989, 0.999) 0.994 (0.988, 0.998) 0.994 (0.984, 0.999)
Tumor Differentiation 0.999 (0.998, 0.999) 0.999 (0.998, 0.999) 0.998 (0.997, 0.998) 0.998 (0.997, 0.998)
PRECURSOR OR BENIGN
Dysplasia 0.947 (0.938, 0.956) 0.949 (0.940, 0.957) 0.931 (0.921, 0.941) 0.912 (0.900, 0.925)
Polyp 0.977 (0.975, 0.980) 0.978 (0.975, 0.980) 0.967 (0.963, 0.970) 0.964 (0.960, 0.967)
No Significant Abnormalities 0.947 (0.944, 0.951) 0.947 (0.943, 0.951) 0.942 (0.938, 0.946) 0.937 (0.933, 0.941)
INFLAMMATORY CONDITION
H. Pylori Helicobacter 0.961 (0.938, 0.979) 0.974 (0.957, 0.987) 0.964 (0.946, 0.979) 0.962 (0.939, 0.980)
Barretts Esophagus 0.988 (0.985, 0.991) 0.990 (0.987, 0.992) 0.989 (0.986, 0.992) 0.962 (0.954, 0.968)
Other Inflammatory Condition 0.957 (0.953, 0.961) 0.957 (0.953, 0.961) 0.945 (0.940, 0.949) 0.962 (0.939, 0.980)
Average 0.977 (0.974, 0.979) 0.978 (0.976, 0.980) 0.971 (0.968, 0.973) 0.965 (0.961, 0.968)
Table A.2.5: Linear probing AUC of detection tasks in the gastrointestinal (GI) tract, spanning cancers, precursors to cancer, inflammatory conditions, and benign conditions. 95% confidence intervals computed with 1000 iterations of bootstrapping are shown in parentheses. The best result per task (p $<$ 0.05) is in bold.
Linear probing PRISM2 PRISM2 PRISM TITAN
Diagnostic Base
Prostate AR 0.912 (0.887, 0.936) 0.917 (0.893, 0.938) 0.897 (0.869, 0.921) 0.911 (0.886, 0.933)
Ovarian FGA 0.877 (0.799, 0.939) 0.891 (0.824, 0.948) 0.847 (0.758, 0.920) 0.859 (0.751, 0.940)
Esophagogastric Her2 0.832 (0.752, 0.902) 0.831 (0.748, 0.905) 0.808 (0.707, 0.897) 0.849 (0.756, 0.925)
Colorectal MSI 0.948 (0.920, 0.970) 0.964 (0.948, 0.979) 0.958 (0.935, 0.977) 0.923 (0.893, 0.951)
Lung EGFR 0.814 (0.783, 0.841) 0.826 (0.799, 0.852) 0.810 (0.782, 0.837) 0.813 (0.784, 0.840)
Melanoma BRAF 0.722 (0.663, 0.784) 0.768 (0.710, 0.823) 0.735 (0.674, 0.791) 0.757 (0.701, 0.814)
Bladder FGFR 0.884 (0.845, 0.920) 0.897 (0.861, 0.927) 0.895 (0.858, 0.929) 0.909 (0.873, 0.943)
Breast CDH1 0.915 (0.889, 0.938) 0.917 (0.890, 0.939) 0.913 (0.888, 0.935) 0.897 (0.870, 0.921)
Endometrial PTEN 0.870 (0.838, 0.900) 0.889 (0.859, 0.917) 0.880 (0.847, 0.908) 0.881 (0.849, 0.909)
Breast PIK3CA 0.674 (0.644, 0.708) 0.677 (0.647, 0.709) 0.686 (0.655, 0.719) 0.658 (0.627, 0.689)
Average 0.845 (0.830, 0.859) 0.857 (0.844, 0.870) 0.843 (0.828, 0.857) 0.846 (0.829, 0.860)
Table A.2.6: Linear probing AUC of biomarker detection per biomarker task. Best results in bold and second best underlined. 95% confidence intervals computed with 1000 bootstrapping iterations are shown in parentheses. With the small data samples, there is not enough statistical power to conclude statistically significant differences; however, the PRISM2 Base embedding tends to outperform all others 7 out of 10 times.
A.3 Zero-shot dialogue completions and contrastive prompts
@@
PanCancer (all and rare)
@@ H. Pylori Helicobacter Is h pylori helicobacter identified in the specimen? {Yes,No}
Table A.3.1: Dialogue completions used for zero-shot evaluation on binary classification tasks. For each task, the language model is prompted with a question. Log probabilities of two selected completions - Yes and No answers - are ranked against each other, with Yes acting as a positive class prediction.
@@
What type of carcinoma is suggested by the specimen, and which subtype is favored?
IDC The specimen contains invasive ductal carcinoma.
ILC The specimen contains invasive lobular carcinoma.
@@
What type of carcinoma is suggested by the specimen, and which subtype is favored?
LUAD The specimen suggests non-small cell carcinoma, with a preference towards adenocarcinoma.
LUSC The specimen suggests non-small cell carcinoma, with a preference towards squamous cell carcinoma.
@@
What type of carcinoma is suggested by the specimen, and which subtype is favored?
CCRCC The specimen reveals renal cell carcinoma of the clear cell type.
PRCC The specimen reveals renal cell carcinoma of the papillary type.
CHRCC The specimen reveals renal cell carcinoma of the chromophobe type.
Table A.3.2: Dialogue completions used for zero-shot evaluation on multi-class classification tasks. For each task, the language model is prompted with a question and a class-specific completion. Log probabilities of the tokens relevant to the class (highlighted in bold) are ranked against each other.
@@
PanCancer (breast subset)
@@
Table A.3.3: Positive- and negative-class prompts used for contrastive zero-shot prediction on the PanCancer dataset. Zero-shot performance is highly sensitive to the prompts used. Using the breast-specific prompts substantially improves PRISM2 contrastive zero-shot performance on the breast subset.
A.4 GPT-4o rewriting and question-answer pair generation prompts
@@
- Never mention physical measurements or quantities using mm or cm or inches, etc.
@@
A.5 Acronyms
@@
LUSC
lung squamous cell carcinoma
LUAD
lung adenocarcinoma
IDC
invasive ductal carcinoma
ILC
invasive lobular carcinoma
DCIS
ductal carcinoma in situ
LCIS
lobular carcinoma in situ

### v2

@@ 4.7 Evaluation datasets
MSK biomarker prediction is an in-house dataset containing 37,633 WSIs gathered from 33,949 samples. The ground truth labels span 10 different biomarkers: Prostate AR, Ovarian FGA, Esophagogastric HER2, Colorectal MSI, Lung EGFR, Melanoma BRAF, Bladder FGFR, Breast CDH1, Endometrial PTEN, and Breast PIK3CA. These labels were originally identified by MSK-IMPACT [10], a targeted test for genetic mutations. The dataset is split into train, tune, and test subsets of size 22,634, 5,606, and 5,709 specimens respectively.
TCGA biomarker prediction is a public dataset containing 10,493 WSIs gathered from 8,611 samples. The ground truth labels are a subset of those in the MSK biomarker prediction dataset: Colorectal MSI, Lung EGFR, Melanoma BRAF, Bladder FGFR, Breast CDH1, Endometrial PTEN, and Breast PIK3CA. Training for each label is done with 3-fold cross-validation.
MSK CRC RFS is an in-house dataset of CRC cases curated for recurrence analysis and RFS prediction. It contains 884 recurrence events across 1,260 cases, spanning 6,513 WSIs. Evaluation on this data is done with 5-fold cross-validation, using month-level precision for the time to event.
TCGA CRC DSS is a set of 13 tasks derived from the public TCGA dataset spanning 9,041 cases with 11,072 WSIs and 2,577 disease-specific death events. DSS is evaluated with 5-fold cross-validation for each of the following tasks: CRC (colorectal cancer: COAD and READ), NSCLC (non-small cell lung carcinoma: LUAD and LUSC), RCC (renal cell carcinoma: KIRK, KIRP, and KICH), BRCA (breast cancer), CNS (central nervous system: GBM and LGG), SARC (sarcoma), ENDO (endometrial: UCEC and UCS), LBTC (liver and biliary tract cancer: LIHC and CHOL), MEL (melanoma: SKCM and UVM), HNSC (head and neck cancer), BLCA (bladder cancer), GI (upper gastrointestinal tract: ESCA and STAD), and PAAD (pancreatic cancer). Tasks were curated such that they have at least 100 DSS events each.
Invasive Breast Cancer Pathology Report Completion is an in-house dataset of 1,638 invasive breast cancer biopsy specimens and their associated pathology reports. The pathology reports are standardized to match the current CAP “Breast Biopsy, Invasive” cancer protocol template. Specimens were selected to ensure the histologic type of the invasive cancer is represented by one of the closed-ended options in the protocol. Macroscopic and quantitative protocol fields were omitted from evaluation. Specifically, procedure type, specimen laterality, tumor site, mitotic rate, and tumor measurement were not included.
Metrics. For linear probing, we use (one-vs-one) area under the curve (AUC) as the primary metric so that model performance could be evaluated without tuning a threshold for the predicted class probabilities. On the other hand, we primarily evaluate direct predictions with balanced accuracy by assuming a threshold at 0.5 probability. For question answering, this is equivalent to taking the model responses directly, without calibration. For fairness, we use product-specific thresholds for clinical-grade products, as these were carefully tuned during product development. Balanced accuracy is the mean recall across all classes which makes it a suitable metric for imbalanced data.
Statistical analysis. Statistical significance was assessed using two-sided paired permutation tests with 1,000 permutations. Confidence intervals of 95% were estimated via bootstrapping using 1,000 iterations. In cases where direct prediction is also reported, metric computation (and thus statistical analyses) are done with the micro-average across folds to enable direct comparison (i.e. no class-reweighting). For TCGA DSS, TCGA biomarkers, and MSK biomarkers the macro-average is used.
5 Acknowledgments
We gratefully thank Philip Rosenfield and Djamilia Dierov for their contributions in making this collaboration possible, Philippe Mathieu for distributed inference support, Mark Fleishman and Sid Senthilnathan for data support, and Wayne Hendricks and Alexander van Eck for infrastructure support.
Appendix S Supplementary Notes
S.1 Direct dialogue completions and contrastive prompts
@@ 5 Acknowledgments
Pan-cancer (all and rare), Prostate / Breast / BLN Product Testing Dataset
@@ H. Pylori Helicobacter Is h pylori helicobacter identified in the specimen? {Yes,No}
Table S1: Yes-no questions used for direct QA-based prediction evaluation on binary classification tasks. For each task, the language model is prompted with a question. Log probabilities of two selected completions - Yes and No answers - are ranked against each other, with Yes acting as a positive class prediction.
@@
What is the histologic type?
IDC A. Invasive ductal carcinoma
ILC B. Invasive lobular carcinoma
@@
What is the histologic type?
LUAD A. Adenocarcinoma
LUSC B. Squamous cell carcinoma
@@
What is the histologic type?
CCRCC A. Clear cell renal cell carcinoma
PRCC B. Papillary renal cell carcinoma
CHRCC C. Chromophobe renal cell carcinoma
Table S2: Multiple-choice questions used for direct QA-based prediction evaluation on multi-class classification tasks. For each task, the language model is prompted with a question and a list of options. Log probabilities of option letters (highlighted in bold) are ranked against each other.
@@
Pan-cancer (breast subset)
@@
Table S3: Positive- and negative-class prompts used for contrastive direct prediction on the Pan-cancer dataset. Contrastive prediction performance is highly sensitive to the prompts used. Using the breast-specific prompts substantially improves PRISM2 contrastive prediction performance on the breast subset.
S.2 GPT-4o rewriting and question-answer pair generation prompts
@@
- Never mention physical measurements or quantities using $m$ or cm or inches, etc.
@@
Multiple-choice question-answer pair generation prompt:
Given a report from routine diagnostic examination of pathology WSIs for a case, create all the multiple choice questions that follow CAP protocols as closely as possible. Exclude measurements, laterality, lymph node counts, and IHC/molecular/cytology information. Do not make unnecessary reference to the history or specimen sections.
Also classify the template that you think best matches the given report according to these categories:
- Breast DCIS, Resection
- Breast DCIS, Biopsy
- Breast Phyllodes Tumor
- Breast Invasive, Resection
- Breast Invasive, Biopsy
- Central Nervous System
- Adrenal Gland
- Appendix NET
- Colon NET
- Duodenum and Ampulla NET
- Jejunum and Ileum NET
- Pancreas (Endocrine)
- Paraganglioma Pheochromocytoma
- Pituitary Neuroendocrine (PitNET)
- Stomach NET
- Thyroid
- Ampulla of Vater
- Anus, Excision
- Anus, Resection
- Appendix
- Colon and Rectum, Biopsy
- Colon and Rectum, Resection
- Distal Extrahepatic Bile Ducts
- Esophagus
- Gallbladder
- GIST, Biopsy
- GIST, Resection
- Hepatocellular Carcinoma
- Intrahepatic Bile Ducts
- Pancreas (Exocrine)
- Perihilar Bile Ducts
- Small Intestine
- Stomach
- Kidney, Biopsy
- Kidney, Resection
- Penis
- Prostate Needle Biopsy Case Level
- Prostate Needle Biopsy Specimen Level
- Prostate, Resection
- Prostate TURP
- Testis Orchiectomy
- Testis Lymphadenectomy
- Ureter, Renal Pelvis, Biopsy
- Ureter, Renal Pelvis, Resection
- Urethra, Biopsy
- Urethra, Resection
- Urinary Bladder, Resection
- Urinary Bladder, Biopsy
- Endometrium Uterus
- Ovary, Fallopian Tube, or Peritoneum
- Trophoblastic Tumors
- Uterine Cervix, Excision
- Uterine Cervix, Resection
- Uterine Sarcoma
- Vagina, Biopsy
- Vagina, Resection
- Vulva
- Larynx
- Oral Cavity
- Major Salivary Glands
- Nasal Cavity and Paranasal Sinuses
- Pharynx
- Cutaneous Squamous Cell Carcinoma
- Precursor and Mature Lymphoid Malignancies
- Myeloid and Mixed / Ambiguous Lineage Neoplasms
- Plasma Cell Malignancies
- Retinoblastoma
- Uveal Melanoma
- Ewing, Resection
- Ewing, Biopsy
- Germ Cell Tumor, Resection
- Germ Cell Tumor, Biopsy
- Hepatoblastoma, Resection
- Hepatoblastoma, Biopsy
- Neuroblastoma, Resection
- Neuroblastoma, Biopsy
- Rhabdomyosarcoma, Resection
- Rhabdomyosarcoma, Biopsy
- Wilms, Resection
- Wilms, Biopsy
- Invasive Melanoma of the Skin: Biopsy
- Invasive Melanoma of the Skin: Excision, Re-Excision
- Merkel Cell Carcinoma
- Lung, Resection
- Diffuse Pleural Mesothelioma
- Thymus
- Bone, Biopsy
- Bone, Resection
- Soft Tissue, Biopsy
- Soft Tissue, Resection
If there is not a matching protocol (i.e. because the sample is benign) then classify as "N/A".
Output your response in a JSON format as shown below:
Example:
user: Endometrioid carcinoma with squamous differentiation is present.
you:
{"cap_protocol": "Endometrium Uterus",
"questions": [{"question": "What is the histologic type?",
"options": [
"Endometrioid carcinoma, NOS",
"Endometrioid carcinoma with squamous differentiation",
"Endometrioid carcinoma, villoglandular variant",
"Endometrioid carcinoma with secretory differentiation",
"Endometrioid carcinoma, other variant (specify): ________________",
"Serous endometrial intraepithelial carcinoma",
"Serous carcinoma",
"Carcinosarcoma (malignant mixed M\"ullerian tumor)",
"Mucinous carcinoma",
"Clear cell carcinoma",
"Small cell neuroendocrine carcinoma",
"Large cell neuroendocrine carcinoma",
"Mixed cell carcinoma (specify types and percentages): __________________________",
"Undifferentiated carcinoma",
"Dedifferentiated carcinoma"
"answer": "Endometrioid carcinoma with squamous differentiation"},
etc.
Make sure that the answer matches one of the options.
S.3 Efficient Implementation
Efficient grouped query attention. We opt to use grouped query attention (GQA) [3] when doing cross attention with tile embeddings. This leads to significant memory savings during training by projecting tile embeddings to a much smaller embedding space as keys/values for cross attention. We use the official FlashAttention [12, 11] kernel from the flash-attn library which has native support for GQA rather than a naïve implementation that repeats key-value heads to match query heads which would negate the potential memory savings. The flash-attn kernel also supports sequence packing which allows us to avoid padding/unpadding before and after attention.
S.4 Acronyms
@@
ADH
atypical ductal hyperplasia
ALH
atypical lobular hyperplasia
AMR
adjusted mean recall
@@
ILC
invasive lobular carcinoma
ITC
infiltrating tumor cell
LCIS
lobular carcinoma in situ
@@
LUAD
lung adenocarcinoma
LUSC
lung squamous cell carcinoma
@@
RCC
renal cell carcinoma
RFS
recurrence-free survival